OpenStudy (anonymous):

Find double integral volume of a solid in first octant bounded by coordinate plane, the cylinder x^2+y^2=4, and z+y=3

OpenStudy (anonymous):

Cylindrical coordinate conversion is your friend here. If \(R\) is the region of interest, then you can describe it by \[R:\left\{(x,y,z)~:~0\le x\le2,~0\le y\le2,~0\le z\le3-y\right\}\] in rectangular coordinates, and \[R:\left\{(r,\theta,z)~:~0\le r\le2,~0\le\theta\le\frac{\pi}{2},~0\le z\le3-r\sin\theta\right\}\] in cylindrical. Does that make sense?

OpenStudy (anonymous):

The volume would then be given by the integral, \[V=\int\int\int_RdV=\int_0^{\pi/2}\int_0^2\int_0^{3-r\sin\theta}r~dz~dr~d\theta\]

OpenStudy (anonymous):

Oh wait, you need a double integral representation... Hmm...

OpenStudy (anonymous):

Thank you so much for the reply. So would V= integral from 0 to 2 integral from 0 to 2 of 3-y dx dy?

OpenStudy (anonymous):

No, you have to convert to polar/cylindrical for this one (double for polar, triple for cylindrical). My definition of \(R\) has a mistake, actually. It should be \[R:\left\{(x,y,z)~:~0\le x\le2,~0\le y\le\color{red}{\sqrt{4-x^2}},~0\le z\le3-y\right\}\] The red is what's preventing you from using rectangular coordinates. Here's the double integral expression: \[V=\int\int_RdV=\int_0^{\pi/2}\int_0^2(3-r\sin\theta)~r~dr~d\theta\] (which is actually equivalent to the triple integral I posted earlier.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw